A ball is projected vertically upwards with an initial velocity of

30m/s. Neglecting air resistance and taking g to be 10m/s2.

a. Find its maximum height

(2)

b. The time taken to return to its starting point.

(3)

c. Plot a velocity-time graph for its downward motion

(4)

d. Predict what would happen to the time of flight if


1
Expert's answer
2022-05-30T15:16:54-0400

"h=\frac {v^2}{2g}=45~m,"

"t=\frac{2v}g=6~s."